OpenStudy (anonymous):

How do i change this? -2

OpenStudy (anonymous):

\[-2\div \sqrt{5} = (-2\sqrt{5})\div5\]

OpenStudy (anonymous):

I don't understand how that is equal.

OpenStudy (anonymous):

multiplying numerator and denominator by \(\sqrt5\) \(\sqrt5\)\(\sqrt5=5\)

OpenStudy (anonymous):

Do you have to multiple -2 too, if not, why?

OpenStudy (anonymous):

-2 remains as -2, from first step to last answer.....

OpenStudy (anonymous):

\(\huge \frac{-2\sqrt5}{\sqrt5\sqrt5}=\frac{-2\sqrt5}{5}\)

OpenStudy (anonymous):

Oh my bad, ty, saw it wrong.

OpenStudy (anonymous):

no problem...